find an equation that satisfies the given conditions
though (-1,-3) perpendicular to the line 2x+7y+9=0

Answer by Alan3354(69443)

find an equation that satisfies the given conditions
though (-1,-3) perpendicular to the line 2x+7y+9=0
---------------
Find the slope of the line
2x+7y+9=0
Solve for y
y = (-2/7)x - 9/7
This is the slope-intercept form y = mx + b
m is the slope = the coefficient of x
m = -2/7
--------
The slope of lines perpendicular is the negative inverse, = +7/2
------
Use m = 7/2 and the point (-1,-3)
y-y1 = m*(x-x1) where (x1,y1) is the point
y+3 = (7/2)*(x+1) is an equation.
------------------
2y+6 = 7x+7
7x - 2y = -1 is another eqn of the same line.